# Erasure-tolerance scheme for the surface codes on neutral atom quantum computers

### v3

@@ Abstract
Neutral atom arrays manipulated with optical tweezers are promising candidates for fault-tolerant quantum computers due to their advantageous properties, such as scalability, long coherence times, and optical accessibility for communication. A significant challenge to overcome is the presence of non-Pauli errors, specifically erasure errors and leakage errors. Previous work has shown that leakage errors can be converted into erasure errors; however, these (converted) erasure errors continuously occur and accumulate over time. Prior proposals have involved transporting atoms directly from a reservoir area—where spare atoms are stored—to the computational area—where computation and error correction are performed—to correct atom loss. While coherent transport is promising, it may not address all challenges—particularly its effectiveness in dense arrays and alternative methods must help. In this study, we evaluate the effects of erasure errors on the surface code using circuit-based Monte Carlo simulations that incorporate depolarizing and accumulated erasure errors. We propose a new scheme to mitigate this problem: a $k$-shift erasure recovery scheme. Our scheme employs code deformation to repeatedly transfer the logical qubit from an imperfect array with accumulated erased qubits to a perfect array, thereby tolerating many accumulated erasures. Furthermore, our scheme corrects erasure errors in the atom arrays while the logical qubits are evacuated from the area being corrected; thus, manipulating optical tweezers for erasure correction does not disturb the qubits that constitute the logical data. Our scheme provides a practical pathway for neutral atom quantum computers to achieve feasible fault tolerance.
@@ I Introduction
Scalable and universal quantum computers are expected to be the next generation beyond classical computers. Among the various platforms, neutral atom quantum computers are particularly promising candidates, as they have demonstrated key functionalities required for fault-tolerant quantum computation [1]. Rydberg atoms possess desirable properties as qubits, such as long coherence times [2, 3, 4] and high controllability in arrays [5, 6, 7].
However, significant hurdles remain to be overcome, including atom loss—the disappearance of qubits caused by the escape of atoms from optical tweezers—and leakage errors, which occur when atom states leak out of the computational qubit space due to imperfect operations during the execution of quantum gates. To address those problems, it has been shown that leakage errors and atom loss are detectable using Rydberg gates [8, 9, 10, 11], and methods to convert leakage errors to erasure errors via Rydberg excitation have also been proposed [12, 13, 14]. Therefore, these techniques allow atom loss and leakage errors to be treated as erasure errors.
@@
A challenge in achieving erasure tolerance with the surface code is refilling vacant spots with reservoir atoms. Previous studies have proposed refilling erased atoms directly into the online array on which an error-correcting code is encoded [17, 12]. Recently, atom transport approaches have been demonstrated, wherein atoms are coherently transported without disturbing surrounding qubits on sparse arrays where the atoms are far apart enough to allow one to move between them [7, 1]. While coherent transport is promising, it may not address all challenges—particularly its effectiveness in dense arrays, which can integrate more qubits within a given space than sparse arrays. Transporting atoms may have side effects that cause atom loss when the transported atoms are close to other atoms, e.g., closer than $5$ $\mu\mathrm{m}$ [5]. This limitation makes the development of alternative methods worthwhile. These methods may complement each other in practical quantum computer architectures by supporting different array densities, operating under varying erasure or physical error rates, and requiring different classical control software.
@@
Initially, we simulate a surface code of distance $d$ with $d$ rounds of measurements, which is the minimum required for the syndrome-measurement circuit with measurement errors. The simulation shows that the surface code does not exhibit a uniform and explicit threshold for the depolarizing error rate and only demonstrates a pseudo-threshold that holds only for relatively short code distances for the depolarizing error rate. In scenarios with a high erasure rate, the surface code at any code distance does not function effectively as an error-correcting code, and no threshold of depolarizing error rate exists. This indicates that accumulated erasure errors can destroy the code. Conversely, in scenarios with a lower erasure error rate, the surface codes function as error-correcting codes and suppress the logical failure rate by increasing the code distance. Consequently, threshold-like behavior can be observed. These results suggest that increasing the code distance at higher erasure error rates makes the code more vulnerable to failure, as the occurrence of erasure errors rises with the number of rounds of syndrome measurements.
@@
To address this problem, we propose a new scheme to tolerate erasure errors, namely the $k$-shift erasure recovery. Our scheme protects the logical state from erasure errors by combining the code deformation technique of the surface code with atom transportation to rearrange or refill atoms in the array. This scheme maintains the array without concern for the decoherence of qubits caused by atom transportation via tweezers, as it separates the operation of maintaining the logical qubit from the maintenance of the atom array. The numerical simulation shows the $k$-shift erasure recovery improves the logical failure rate compared with the simple repetition of syndrome measurements on a single array.
@@
II Rydberg Atom Qubits in 2-D Arrays and Their Error Model
@@
An optical tweezer is a technique used to trap a single atom at a spot focused to micrometer-scale widths, shaped by an objective lens and a spatial light modulator (SLM) or an acousto-optic deflector (AOD)[10]. Single-atom arrays have the capability for dynamic reconfiguration and rearrangement using these optical techniques[6, 7].
II.1 Rydberg Atom Qubits
A single atom has multiple energy levels; two of these levels can be chosen to define a qubit, such as levels within a hyperfine structure. Arbitrary single-qubit gates are realized using Rabi oscillations between these qubit sublevels [10]. Multi-qubit gates are achieved via dipole-dipole interactions between atoms in Rydberg states, known as Rydberg gates.
Let $r$ be the distance between atoms. The dipole-dipole interaction decreases proportionally to $1/r^{6}$. Rydberg gates can be applied between atoms that are not necessarily nearest neighbors in the array, provided they are close enough for the dipole interaction to act effectively. The typical two-qubit Rydberg gate is the controlled-Z (CZ) gate [18]. A controlled-NOT (CNOT) gate can be constructed by surrounding a CZ gate with Hadamard gates on the target qubit. By combining arbitrary single-qubit gates and multi-qubit gates, a universal quantum gate set is achieved on neutral atom quantum computers.
II.2 Leakage Errors, Atom Loss in Neutral Atom Qubits, and Erasure Errors
Neutral atom quantum computers with optical tweezer arrays are subject not only to Pauli errors within the qubit space but also to leakage errors, where the state leaks to levels outside of the qubit space, and atom loss, where atoms disappear from optical tweezer sites. Some leakage states can be returned to the qubit space through optical pumping [17], while other leakage errors can be converted to erasure errors if their occurrence is detected [12]. Two-electron neutral atoms have demonstrated the erasure conversion protocol for leaked states due to radiative decay and blackbody radiation errors [13, 14].
@@ II.2 Leakage Errors, Atom Loss in Neutral Atom Qubits, and Erasure Errors
$\mathcal{E}_{e}(\rho)=(1-p_{\rm e})\rho+p_{\rm e}\ket{e}\bra{e},$ (1)
where |$\ket{e}$ is a state outside the qubit space [19].
@@ III.1 Surface Code
where $v$, $e$, and $p$ denote the vertices, edges, and faces of the square lattice, respectively. Here, $\partial$ and $\partial^{*}$ are the boundary operators of the square lattice and its dual lattice. The boundaries of the code are classified into two types: the smooth boundary, where $X_{s}$ terminates, and the rough boundary, where $Z_{p}$ terminates.
@@
Figure 1: (a) Illustration of the stabilizer generators $\mathcal{\tilde{G}}$ of the surface code. The large white circles represent data qubits. The orange (blue) squares represent stabilizer generators of $XXXX$ ($ZZZZ$) for star (plaquette) operators. The small orange (blue) circles are ancilla qubits for syndrome measurements. Suppose a data qubit indicated by a circle with diagonal lines disappears; in that case, the stabilizer generators around it are merged. Then, a super star (plaquette) operator of a large orange (blue) polygon is added as a new stabilizer generator to maintain the functionality as an error-correcting code. (b) An example of the time evolution of a qubit array encoded by the surface code with distance $5$. The bottom layer illustrates the situation before the syndrome measurement, with no erased qubits. The vertical axis represents time, showing how the erased qubits (circles with diagonal lines) increase as time progresses. (c) A circuit diagram of syndrome measurements for the surface code. The shaded regions in blue and orange represent syndrome measurements of $X^{\otimes 4}$ and $Z^{\otimes 4}$, respectively.
@@ III.2 Error Detection and Decoding
The encoded state |$\ket{\psi_{L}}$ is stabilized by the group $\mathcal{\mathcal{S}}$ as constructed above. Errors in this state are detected by measuring the eigenvalues corresponding to the stabilizer generators $S\in\mathcal{G}$, yielding either $+1$ or $-1$, which provides information about the Pauli errors that have occurred in the qubit array. This information is obtained from indirect measurements of the star (plaquette) operators as shown in Fig. 1c using the orange (blue) ancilla qubits in Fig. 1a. This process is called the syndrome measurement.
@@
The error syndromes contain information about the errors. A matching graph $\mathcal{N}$, in which vertices represent spatiotemporal flips of error syndromes from $+1$ to $-1$ or vice versa [22, 23, 24, 25], is formed to decode these syndromes into Pauli errors. The detector error model is a typical method to form $\mathcal{N}$ from syndrome measurement circuits [26]. Then, a Minimum Weight Perfect Matching (MWPM) problem solver [27, 25] matches the error-detected vertices to find the most likely error chains.
@@ III.3 Previous Work on Erasure Errors
III.3.1 Static Erasure Errors
In systems where erasure errors occur statically, the positions and number of erasure errors are known before executing quantum computation, and they remain unchanged during execution. A typical case of static erasure errors is fabrication defects in superconducting quantum circuits [16, 28]. Although static erasure errors do not emerge during computation but are present beforehand, they share many characteristics with dynamic erasure errors: they are caused by the absence of qubits, their locations are known (heralded), and stabilizer deformation is required to deal with them, as explained in the following sections.
This type of erasure can be corrected using super stabilizers. There are two main methods for forming super stabilizers: the method of Stace and Barrett [15] and the method of Nagayama et al. [16]. The first method is simple and easy to implement; super stabilizers are created by multiplying stabilizers associated with the erased data qubits, as illustrated in Fig. 1a. In this method, star and plaquette stabilizers are merged into super stabilizers symmetrically. However, this method generates junk qubits as side effects—these are extra degrees of freedom arising from the reduction of two stabilizers due to an erased data qubit. Although junk qubits do not cause logical errors, they can decrease the threshold of the physical error rate. The second method is an improved version that uses both super stabilizers and weight-reduced stabilizers. This method avoids generating junk qubits but results in an asymmetric deployment of super stabilizers.
III.3.2 Instantly Correctable Erasure Errors
Some systems may experience erasure errors dynamically during computation, where the number of erasure errors changes over time. Mechanisms for detecting the positions of erased qubits have been investigated, depending on physical and architectural systems [29, 17, 12, 11].
There are two types of such dynamic erasure errors. Erasure errors of the first type remain unless corrected. A typical example of this type is neutral atom quantum computers; erasure errors remain unless special operations are executed to fix them, such as refilling the erased position with a reservoir atom instantly. This refilling is achieved by transporting atoms using optical tweezers. In the erasure conversion approach [12, 30], the original qubit information is exactly erased, but the atom may remain in its position. In this case, the erased qubits, which are in a completely mixed state, can be incorporated into the surrounding stabilizers and projected onto the proper codeword state of the quantum error-correcting code.
Erasure errors of the second type do not persist. A typical example of this type is one-way quantum computation using a 3-D cluster state [29, 31, 30]. In this model, all qubits are eventually measured and disrupted, so erasure errors in a layer corresponding to a syndrome measurement round do not persist into the next round. This model can tolerate an erasure rate of up to $25\%$ [31], even though we do not actually “correct” the erased qubits.
III.3.3 Accumulating Dynamic Erasure Errors
Some systems cannot correct dynamic erasure sequentially after detecting their occurrence [32]. In dense arrays of neutral atom quantum computers, sequential erasure correction will be difficult because transporting atoms into the online array can disturb other qubits in the case of dense arrays [5]. Consequently, the number of erasure errors in the array increases over time. In such cases, the surface code does not function properly because the distribution of erased qubits changes continuously, and error detection and decoding do not work effectively around erased qubits.
@@ IV.1 Main Problem
We consider a situation where erasure errors occur dynamically and cannot be corrected immediately after detection. Neutral atom quantum computers with optical tweezer arrays, the dense one, correspond to this scenario since atoms cannot be transported to the dense online array on which logical qubits are encoded because atom transport possibly causes adverse effects such as decoherence or repulsion between atoms.
@@ IV.3 Composition of Super Stabilizers and Decoding under Dynamic Erasure
Then, the matching graph $\mathcal{N}$ is formed from the syndromes that are measured with the updated stabilizer generators. Finally, an MWPM solver decodes the syndromes, and we achieve Pauli error correction under the presence of erasure errors.
@@
1:function union_find($stabilizers$)
@@
6:              if $qubit\;is\;erased$ then
@@
14:function create_stabilizers($ideal\_stabilizers$)
@@
17:     $new\_stabilizers\leftarrow\textsc{merge\_stabilizers}(root,ideal\_stabilizers)$triangleright This function merges stabilizers belonging to the same tree using Stace’s method.
18:     return $new\_stabilizers$

### v4

@@ Abstract
Neutral atom arrays manipulated with optical tweezers are promising candidates for fault-tolerant quantum computers due to their advantageous properties, such as scalability, long coherence times, and optical accessibility for communication. A significant challenge to overcome is the presence of non-Pauli errors, specifically erasure errors and leakage errors. Previous work has shown that leakage errors can be converted into erasure errors; however, these (converted) erasure errors continuously occur and accumulate over time. Prior proposals have involved transporting atoms directly from a reservoir area—where spare atoms are stored—to the computational area—where computation and error correction are performed—to correct atom loss. While coherent transport is promising, it may not address all challenges—particularly its effectiveness in dense arrays and the efficacy of alternative methods should be investigated. In this study, we evaluate the effects of erasure errors on the surface code using circuit-based Monte Carlo simulations that incorporate depolarizing and accumulated erasure errors. We propose a new scheme to mitigate this problem: a $k$-shift erasure recovery scheme. Our scheme employs code deformation to repeatedly transfer the logical qubit from an imperfect array with accumulated erased qubits to a perfect array, thereby tolerating many accumulated erasures. Furthermore, our scheme corrects erasure errors in the atom arrays while the logical qubits are evacuated from the area being corrected; thus, manipulating optical tweezers for erasure correction does not disturb the qubits that constitute the logical data. Our scheme provides a practical pathway for neutral atom quantum computers to achieve feasible fault tolerance.
@@ I Introduction
Scalable and universal quantum computers are expected to be the next generation beyond classical computers. Among the various platforms, neutral atom quantum computers are particularly promising candidates, as they have demonstrated key functionalities required for fault-tolerant quantum computation [1]. Rydberg atoms possess desirable properties as qubits, such as long coherence times [2, 3, 4, 5] and high controllability in arrays [6, 7, 8, 5].
However, significant hurdles remain to be overcome, including atom loss—the disappearance of qubits caused by the escape of atoms from optical tweezers—and leakage errors, which occur when atom states leak out of the computational qubit space due to imperfect operations during the execution of quantum gates. To address those problems, it has been shown that leakage errors and atom loss are detectable using Rydberg gates [9, 10, 11, 12], and methods to convert leakage errors induced by Rydberg excitation to erasure errors have also been proposed [13, 14, 15]. Therefore, these techniques allow atom loss and leakage errors to be treated as erasure errors.
@@
A challenge in achieving erasure tolerance with the surface code is refilling vacant spots with reservoir atoms. Previous studies have proposed refilling erased atoms directly into the online array on which an error-correcting code is encoded [18, 13]. Recently, atom transport approaches have been demonstrated, wherein atoms are coherently transported without disturbing surrounding qubits on sparse arrays where the atoms are far apart enough to allow one to move between them [8, 1]. While coherent transport is promising, it may not address all challenges—particularly its effectiveness in dense arrays, which can integrate more qubits within a given space than sparse arrays. Transporting atoms may have side effects that cause atom loss when the transported atoms are close to other atoms, e.g., closer than 55 $\mu\mathrm{m}$ [6]. This limitation makes the development of alternative methods worthwhile. These methods may complement each other in practical quantum computer architectures by supporting different array densities, operating under varying erasure or physical error rates, and requiring different classical control software.
@@
Initially, we simulate a surface code of distance $d$ with $d$ rounds of measurements, which is the minimum required for the syndrome-measurement circuit with measurement errors. The simulation shows that the surface code does not exhibit a uniform and explicit threshold for the depolarizing error rate and only demonstrates a pseudo-threshold that holds only for relatively short code distances for the depolarizing error rate under dynamically accumulating erasures. In scenarios beyond a certain erasure rate, the surface code at any code distance does not function effectively as an error-correcting code, and no threshold of depolarizing error rate exists. This indicates that accumulated erasure errors can destroy the code. Conversely, in scenarios with a lower erasure error rate, the surface codes function as error-correcting codes and suppress the logical failure rate by increasing the code distance. Consequently, threshold-like behavior can be observed. These results suggest that increasing the code distance at higher erasure error rates makes the code more vulnerable to failure, as the occurrence of erasure errors rises with the number of rounds of syndrome measurements.
@@
To address this problem, we propose a new scheme to tolerate erasure errors, namely the $k$-shift erasure recovery. Our scheme protects the logical state from erasure errors by combining the code deformation technique of the surface code with atom transportation to rearrange or refill atoms in the array. This scheme maintains the array without concern for the atom loss caused by atom transportation via tweezers, as it separates the operation of maintaining the logical qubit from the maintenance of the atom array. The numerical simulation shows the $k$-shift erasure recovery improves the logical failure rate compared with the simple repetition of syndrome measurements on a single array.
@@
II Rydberg Atom Qubits in Optical Tweezer Arrays and Their Error Model
@@
Optical tweezer arrays are arrays of tightly focused laser spots, generated by sending laser beams into a high-NA objective lens. Two common techniques used to generate such arrays are spatial light modulators (SLMs) and acousto-optic deflectors (AODs) [11]. SLMs can create arbitrarily-shaped static tweezer arrays, while AODs specialize in moving atoms on a microsecond timescale. This combination is highly favorable for realizing large, reconfigurable atom arrays [20, 7, 8].
A single atom has multiple energy levels; two of these levels can be chosen to define a qubit, such as levels within a hyperfine structure. Arbitrary single-qubit gates are realized using Rabi oscillations between these qubit sublevels [11]. Multi-qubit gates are achieved via van der Waals interactions between atoms in Rydberg states, known as Rydberg gates.
Let $r$ be the distance between atoms. The dipole-dipole interaction decreases proportionally to $1/r^{6}$. Rydberg gates can be applied between atoms that are not necessarily nearest neighbors in the array, provided they are close enough for the dipole interaction to act effectively [21, 22]. The typical two-qubit Rydberg gate is the controlled-Z (CZ) gate [19]. A controlled-NOT (CNOT) gate can be constructed by surrounding a CZ gate with Hadamard gates on the target qubit. By combining arbitrary single-qubit gates and multi-qubit gates, a universal quantum gate set is achieved on neutral atom quantum computers.
Neutral atom qubits on the tweezer arrays are subject not only to Pauli errors within the qubit space but also to leakage errors, where the state leaks to levels outside of the qubit space, and atom loss, where atoms disappear from optical tweezer sites. Some leakage states can be returned to the qubit space through optical pumping [18], while other leakage errors can be converted to erasure errors if their occurrence is detected [13]. Alkaline-earth atoms have demonstrated the erasure conversion protocol for leaked states due to Raman scattering from the tweezer light within the qubit manifold [14] and imperfect state preparation [15].
@@
$|,\mathcal{E}_{e}(\rho)=(1-p_{\rm e})\rho+p_{\rm e}\ket{e}\bra{e}$, (1)
[…]
The erasure correction for these two types of erasure errors are assumed to be achieved by atom replacement in this work. For the atom loss, we refer to the process of reintroducing an atom into a tweezer site where it has been physically atom lost as atom replacement. For the leakage error, we assume that the state cannot be returned to the computational basis when a leakage error is detected. In such cases we physically remove atoms under leakage error and correct the vacant tweezer sites using atom replacement. Therefore, the both erasure corrections are achieved by atom replacement.
@@ III.1 Surface Code
where $v$, $e$, and $p$ denote the vertices, edges, and faces of the square lattice, respectively. Here, $\partial$ and $\partial^{*}$ are the boundary operators of the square lattice and its dual lattice. The boundaries of the code are classified into two types: the smooth boundary, where $X_{s}$ appears at the edge, and the rough boundary, where $Z_{p}$ appears at the edge.
@@
Figure 1: (a) Illustration of the stabilizer generators $\mathcal{\tilde{G}}$ of the surface code. The large white circles represent data qubits. The orange (blue) squares represent stabilizer generators of XXXXXXXX (ZZZZZZZZ) for star (plaquette) operators. The small orange (blue) circles are ancilla qubits for syndrome measurements. Suppose a data qubit indicated by a circle with diagonal lines disappears; in that case, the stabilizer generators around it are merged. Then, a super star (plaquette) operator of a large orange (blue) polygon is added as a new stabilizer generator to maintain the functionality as an error-correcting code. (b) An example of the time evolution of a qubit array encoded by the surface code with distance 55. The bottom layer illustrates the situation before the syndrome measurement, with no erased qubits. The vertical axis represents time, showing how the erased qubits (circles with diagonal lines) increase as time progresses. (c) A circuit diagram of syndrome measurements for the surface code. The shaded regions in blue and orange represent syndrome measurements of $X^{\otimes 4}$ and $Z^{\otimes 4}$, respectively.
@@ III.2 Error Detection and Decoding
[…]
@@
The error syndromes contain information about the errors. A matching graph $\mathcal{N}$, in which vertices represent spatiotemporal flips of error syndromes from +1+1 to 1-1 or vice versa [26, 27, 28, 29], is formed to decode these syndromes into Pauli errors. The detector error model is a typical method to form $\mathcal{N}$ from syndrome measurement circuits [30]. Then, a Minimum Weight Perfect Matching (MWPM) problem solver [31, 29] matches the error-detected vertices to find the most likely error chains.
@@ III.3 Previous Work on Erasure Errors
In systems where erasure errors occur statically, the positions and number of erasure errors are known before executing quantum computation, and they remain unchanged during execution. A typical case of static erasure errors is fabrication defects in superconducting quantum circuits [17, 32]. This type of erasure can be corrected using super stabilizers. There are two main methods for forming super stabilizers: the method of Stace and Barrett [16] and the method of Nagayama et al. [17]. The first method is simple and easy to implement; super stabilizers are created by multiplying stabilizers associated with the erased data qubits, as illustrated in Fig. 1a.
Some systems may experience erasure errors dynamically during computation, where the number of erasure errors changes over time. Mechanisms for detecting the positions of erased qubits have been investigated, depending on physical and architectural systems [33, 18, 13, 12]. There are two types of such dynamic erasure errors. Erasure errors of the first type remain unless corrected. A typical example of this type is neutral atom quantum computers; erasure errors remain unless special operations are executed to fix them, such as refilling the erased position with a reservoir atom instantly. This refilling is achieved by transporting atoms using optical tweezers. In the erasure conversion approach [13, 34], the original qubit information is exactly erased, but the atom may remain in its position. In this case, the erased qubits, which are in a completely mixed state, can be incorporated into the surrounding stabilizers and projected onto the proper codeword state of the quantum error-correcting code.
Some systems cannot correct dynamic erasure sequentially after detecting their occurrence [35]. In dense arrays of neutral atom quantum computers, sequential erasure correction will be difficult because transporting atoms into the online array can disturb other qubits [6]. Consequently, the number of erasure errors in the array increases over time. In such cases, the surface code does not function properly because the distribution of erased qubits changes continuously, and error detection and decoding do not work effectively around erased qubits.
@@ IV.1 Main Problem
We consider a situation where erasure errors occur dynamically and cannot be corrected immediately after detection. Neutral atom quantum computers with dense optical tweezer arrays correspond to this scenario since atoms cannot be transported to the dense online array on which logical qubits are encoded because atom transport possibly causes adverse effects such as decoherence or repulsion between atoms.
@@ IV.3 Composition of Super Stabilizers and Decoding under Dynamic Erasure
Then, the matching graph $\mathcal{N}$ is formed from the syndromes that are measured with the updated stabilizer generators. Finally, an MWPM solver decodes the syndromes, and we achieve Pauli error correction in the presence of erasure errors.
@@
1:function union_find(stabilizersstabilizers)
@@
6:     if qubitiserasedqubit\;is\;erased then
@@
14:function create_stabilizers(ideal_stabilizersideal\_stabilizers)
@@
17:  $new\_stabilizers\leftarrow\textsc{merge\_stabilizers}(root,ideal\_stabilizers)$ $\triangleright$ This function merges stabilizers belonging to the same tree using Stace’s method.
18:  return new_stabilizersnew\_stabilizers